a wire of resistance 5 ohms is bent to form a square what is the equal and resistant in ohm across the diagonally opposite vertices of the square

Grade:10

2 Answers

jitender
6 years ago
R~lResistance directly proportional to the length of the site.. Let total length be 4lThen resistance of each side of square is r/4Now along the diagonal...Two sides are in series and are in parralel with the other two.So Rs=r/2And Rp=r/4So r across diagonally opposite ends is R/4
Shreyesh gunda
6 years ago
R is the answer.If the four wires of each resistnace is R and attached to form a square .then find the rrsistance between the vertices of a square.